OpenStudy (anonymous):

Give the vector equation of the line passing through P = (1,-2) and Q = (3,0)

OpenStudy (arnavguddu):

R = OA + t(AB) OA = vector of 1st point say = (1,-2) AB = vector of the difference between the 2 given points in ur case (3-1,0+2) = (2,2) t is just any value